# Asymmetric Bias in Text-to-Image Generation with Adversarial Attacks

### 22 Dec 2023 v1

@@ C.2 Score Function
Here $w_{t}$ and $w_{s}$ are weighing scalars and cos denotes the standard cosine similarity metric between two one-dimensional text embeddings. For simplicity, we set $w_{t}=w_{s}=1$ for all experiments.
C.3 Optimization over Discrete Tokens
The main challenge in optimizing $\mathcal{S}$ is that we have to optimize over a discrete set of tokens. Moreover, since the search space is exponential ($k^{|V|}$ for k suffix tokens), a simple greedy search is intractable. However, we can leverage gradients with respect to the one-hot tokens to find a set of promising candidates for replacement at each token position. We use the negated Score Function as the loss function $\mathcal{L}(x_{1:n})=-\mathcal{S}(x_{1:n})$. Maximizing the score is equivalent to minimizing the loss. Since losses are used for top K token selection, the absolute value of the loss doesn’t matter. We can compute the linearized approximation of replacing the $i^{t}h$ token i, $x_{i}$ by evaluating the gradient
@@ C.3 Optimization over Discrete Tokens
Where ${e_{x_{i}}}$ denotes the one-hot vector representing the current value of the $i^{t}h$ token. Taking gradient with respect to one-hot vectors was pioneered by HotFlip Ebrahimi et al. (2017) and applied on Stable Diffusion by a concurrent work Yang et al. (2023a). Based on this heuristic, we presented two algorithms for finding adversarial suffix tokens against Stable Diffusion.
Appendix D Additional Determinants of Attack Success
These sections discuss the factors beyond the asymmetric properties that are tied to the attack success rate. We have found factors like whether target token synonyms are allowed, attack suffix length and attack POS types are factors indicating the attack’s success. We also found unlike LLM attacks, adversarial suffixies do not transfer across T2I indicating these models might be harder to attack than single modality models.
D.1 Restricted Token Selection
Emulating QFAttack.
@@ D.1 Restricted Token Selection
Blocking Selection of Target Tokens.
Another potential use case is preventing the selection of the exact target word. However, we find that the algorithm simply finds a synonym or sub-word tokenization for the target word when the exact target word (token) is restricted. For example, when attempting to attack the input text “a backpack on a mountain.” to “a castle on a mountain.”, restricting the token corresponding to “castle” leads to the algorithm including synonyms like “palace”, “chateau”, “fort” or sub-word tokenization like “cast le” or “ca st le” in the adversarial suffix. We find that the effectiveness of the algorithm isn’t affected when the exact target token is restricted and it still finds successful adversarial suffixes using synonyms (when preconditions are met).
@@
We set the number of adversarial tokens to $k=5$ for all experiments. However, we observe that not all input text-target text pairs require $k=5$. “a red panda/car in a forest.” can be attacked with a few as $k=2$, i.e. “a red panda/car in a forest.” while “a guitar/piano in a music store.” required all $k=5$ (see Appendix B). We leave the comprehensive study on the effect of changing the number of tokens as future work.
D.2 Certain Adjectives Resist Adversarial Attacks
We observed that adversarial attacks targeting certain adjectives like color had a very low ASR. For example, swapping out “red” with “blue” in the prompt “a red car in a city road.” failed in all instances. Further challenging examples include “a red/purple backpack on a mountain.” and “a white/black swan on a lake.”. However, other adjectives like “a sapling/towering tree in a forest” or “a roaring/sleeping lion in the Savannah.” had high ASR in at least one direction. We leave further analysis of this phenomenon as future work.
D.3 Adversarial Suffixes Do Not Transfer across T2I Models
@@
Appendix E Human Evaluation WebUI

### 14 Feb 2024 v2

@@ E.2 Score Function
Here, $w_{t}$ and $w_{s}$ are weighing scalars and cos denotes the standard cosine similarity metric between two one-dimensional text embeddings. For simplicity, we set $w_{t}=w_{s}=1$ for all experiments.
E.3 Optimization over Discrete Tokens
The main challenge in optimizing $\mathcal{S}$ is that we have to optimize over a discrete set of tokens. Furthermore, since the search space is exponential ($k^{|V|}$ for k suffix tokens), a simple greedy search is intractable. However, we can leverage gradients with respect to the one-hot tokens to find a set of promising candidates for replacement at each token position. We use the negated Score Function as the loss function $\mathcal{L}(x_{1:n})=-\mathcal{S}(x_{1:n})$. Maximizing the score is equivalent to minimizing the loss. Since losses are used for top K token selection, the absolute value of the loss does not matter. We can compute the linearized approximation of replacing the $i^{t}h$ token i, $x_{i}$ by evaluating the gradient
@@ E.3 Optimization over Discrete Tokens
Here ${e_{x_{i}}}$ denotes the one-hot vector that represents the current value of the $i^{t}h$ token. Taking gradient with respect to one-hot vectors was pioneered by HotFlip Ebrahimi et al. (2017) and applied on Stable Diffusion by a concurrent work Yang et al. (2023a). Based on this heuristic, we presented two algorithms for finding adversarial suffix tokens against Stable Diffusion.
Appendix F Primary Determinants of Attack Success
(a) “a sofa and a bed in a room.”
(b) “a dragon guarding a treasure.”
Figure 8: Examples of prompts that have low Base Success Rate (BSR) that highlight cases where Stable Diffusion fails to generate images that match the input prompt.
(a) ASR vs. Perplexity Difference ($\Delta_{1}$ in Eqn. 6)
(b) ASR vs. BSR (of target text)
(c) ASR vs. Baseline Distance Difference ($\Delta_{2}$ in Eqn. 7)
(d) ASR for Negative and Positive $\Delta_{2}$
Figure 9: Correlation of ASR on $\Delta_{1}$, $\Delta_{2}$ and BSR. Data is reported using the Multiple Token Perturbation algorithm on HQ-Pairs. We find that the Perplexity Difference $\Delta_{1}$ does not correlate with ASR. BSR shows a weak positive correlation and Baseline Distance Difference $\Delta_{2}$ shows a moderate negative correlation with ASR.
Appendix G Additional Determinants of Attack Success
These sections discuss factors beyond the asymmetric properties that are related to the success rate of the attack. We have found factors like whether target token synonyms are allowed, attack suffix length and attack POS types are factors indicating the attack’s success. We also found that, unlike LLM attacks, adversarial suffixes do not transfer across T2I, indicating that these models might be harder to attack than single-modality models.
G.1 Restricted Token Selection
Emulating QFAttack
@@ G.1 Restricted Token Selection
Blocking Selection of Target Tokens
Another potential use case is preventing the selection of the exact target word. However, we find that the algorithm simply finds a synonym or subword tokenization for the target word when the exact target word (token) is restricted. For example, when attempting to attack the input text “a backpack on a mountain.” to “a castle on a mountain.”, restricting the token corresponding to “castle” leads to the algorithm including synonyms like “palace”, “chateau”, “fort” or subword tokenization like “cast le” or “ca st le” in the adversarial suffix. We find that the effectiveness of the algorithm isn’t affected when the exact target token is restricted and it still finds successful adversarial suffixes using synonyms (when preconditions are met).
@@
We set the number of adversarial tokens to $k=5$ for all experiments. However, we observe that not all input text-target text pairs require $k=5$. “a red panda/car in a forest.” can be attacked with a few as $k=2$, i.e. “a red panda/car in a forest.” while “a guitar/piano in a music store.” required all $k=5$ (see Appendix D). We leave a comprehensive study on the effect of changing the number of tokens for future work.
G.2 Certain Adjectives Resist Adversarial Attacks
We observed that adversarial attacks targeting certain adjectives, such as color, had a very low ASR. For example, swapping out “red” with “blue” in the prompt “a red car on a city road.” failed in all instances. Further challenging examples include “a red/purple backpack on a mountain.” and “a white/black swan on a lake.”. However, other adjectives like “a sapling/towering tree in a forest” or “a roaring/sleeping lion in the Savannah.” had high ASR in at least one direction. We leave further analysis of this phenomenon for future work.
G.3 Adversarial Suffixes Do Not Transfer across T2I Models
@@
Appendix H Human Evaluation WebUI